# A Survey on Optimization Studies of Group Centrality Metrics

### v1

@@ 3 Brief history on group centrality
Centrality and of graphs and classes (Everett and Borgatti, 1999) Extending centrality (Everett and Borgatti, 2005) Identifying sets of key players in a social network (Borgatti, 2006) Finding the most prominent group in complex networks (Puzis et al., 2007) On variants of shortest-path betweenness centrality and their generic computation (Brandes, 2008) Incremental deployment of network monitors based on group betweenness centrality (Dolev et al., 2009) Routing betweenness centrality (Dolev et al., 2010) An integer programming approach for finding the most and the least central cliques (Vogiatzis et al., 2015) Efficient maximum closeness centrality group identification (Chen et al., 2016) pt Finding groups with maximum betweenness centrality (Veremyev et al., 2017) Maximizing network performance based on group centrality by creating most effective k-links (Ohara et al., 2017) I/O-efficient calculation of H-group closeness centrality over disk-resident graphs (Zhao et al., 2017) pt Finding clique clusters with the highest betweenness centrality (Rysz et al., 2018) On maximum degree-based $\gamma$-quasi-clique problem: Complexity and exact approaches (Pastukhov et al., 2018) pt A new group centrality measure for maximizing the connectedness of network under uncertain connectivity(Fushimi et al., 2019) Identification of essential proteins using induced stars in protein–protein interaction networks (Vogiatzis and Camur, 2019) Finding critical links for closeness centrality (Veremyev et al., 2019) Current flow group closeness centrality for complex networks? (Li et al., 2019) pt Maximum closeness centrality k-clubs: a study of dock-less bike sharing (Rahim Taleqani et al., 2020) Detecting a most closeness-central clique in complex networks (Nasirian et al., 2020) Group centrality algorithms based on the h-index for identifying influential nodes in large-scale networks (Li et al., 2020) k-step betweenness centrality (Akgün and Tural, 2020) Group centrality maximization for large-scale graphs (Angriman et al., 2020) Finding influential groups in networked systems: The most degree-central clique problem (Zhong et al., 2021) pt The star degree centrality problem: A decomposition approach (Camur et al., 2022) Novel centrality metrics for studying essentiality in protein-protein interaction networks based on group structures (Rasti and Vogiatzis, 2022) PRESTO: Fast and Effective Group Closeness Maximization(Rajbhandari et al., 2022) pt Finding groups with maximum betweenness centrality via integer programming with random path sampling (Lagos et al., 2023) Finding the most degree-central walks and paths in a graph: Exact and heuristic approaches (Matsypura et al., 2023) The stochastic pseudo-star degree centrality problem (Camur et al., 2023) Exact Algorithms for Group Closeness Centrality (Staus et al., 2023)
@@
Figure 4: The emergence of OR scientists during the evolution of studies regarding group centrality metrics in the last two decades.
@@
Pioneering the field are Everett and Borgatti (Everett and Borgatti, 1999) who are the first in the OR literature to discuss extending centrality and importance to groups of nodes in a network. They motivate the question for groups that are known a priori. We quote from their work: “Are the lawyers more central than the accountants in a given organisation’s social network? Is one particular ethnic minority more integrated into the community than another? To what extent are particular groups or classes (women, the elderly, African-Americans, etc.) marginalized in different networks?”. This a priori availability of group information that can help us derive descriptive information about the group itself. However, as the authors also note in that same article, we may also focus on identifying groups based on the structure of the network itself (Everett and Borgatti, 1999). They do mention cliques as a motivation for clique centrality; see the later work by Vogiatzis et al. (Vogiatzis et al., 2015).
@@
In their original (Everett and Borgatti, 1999) and subsequent work (Everett and Borgatti, 2005), they base their proposed metrics to nodal centrality metrics. Specifically, they extend traditional metrics in the form of degree, closeness, betweenness, and flow centrality to accommodate a group setting. The authors and subsequent studies have shown how individual nodal calculations for centrality do not necessarily capture the calculations that are necessary for group centrality. In the early 2000s, a fundamental study by Newman and Girvan (Newman and Girvan, 2004) (not included in the diagram of Figure 4) use betweenness centrality to identify groups within a network. They do so by calculating edge betweenness and removing the edge of highest score until the network is disconnected. Importantly, this metric is recalculated after ever removal. As this work has an arguably different focus than our review, we do not include it in our history. For completeness, though, and as readers may want to investigate the intersection between community detection and centrality, we offer the interesting comparison made by Danon et al. in (Danon et al., 2005).
@@
Continuing with these earlier years of group importance, in (Borgatti, 2006) Borgatti utilizes group centrality in social network analysis to solve a key player problem. The key player problem (also known as the critical node detection problem) tasks itself with identifying a set of nodes that are “critical” for the well-being or success of a network. More recent extensions have since focused on identifying sets of network elements (nodes, edges, cliques). Borgatti (2006) proposes one of the very first combinatorial optimization approaches in this domain. To our best knowledge, this is one of the earliest studies where optimization algorithms (in the form of genetic algorithms) are employed in the context of group centrality. The importance of players are measured through the impact of their absence on the graph cohesiveness. The author also discusses how degree, betweenness, and closeness metrics may fall short when applied “as is” for this problem definition of key players. Subsequently, in Puzis et al. (2007) the authors study the problem of determining groups with the largest betweenness in a large-scale network. They apply their algorithms on a network service provider infrastructure data set to optimally distribute intrusion detection devices.
@@
Another prominent work by Brandes (Brandes, 2008) presents a wide range of variants of betweenness centrality including proximal betweenness, bounded-distance betweenness, distance-scaled betweenness, and edge betweenness. That same work also presents an algorithm to compute the group betweenness of a “given” group. It is critical to mention that the paper focuses on computing the centrality of a fixed group. This is an important step, and one that subsequent studies have used. In this work, the high complexity associated with identifying the optimal group in terms of group betweenness is also mentioned as an important factor. Thus, the study lacks algorithmic details in detecting the most betweenness group, different than Puzis et al. (2007).
@@
Nonetheless, Dolev, Elovici, and Puzis Dolev et al. (2009) further this research by studying group betweenness centrality measured as the information flow in an evolving network. The authors propose a heuristic approach with an approximation guarantee that generalizes the heuristic algorithm proposed in Puzis et al. (2007) for a similar problem setting. Later, Fink and Spoerhase (2011) improves the analysis of Dolev et al. (2009). However, since none of these algorithms scale well in complex networks, in Mahmoody et al. (2016) the authors design a randomized algorithm guaranteeing an approximation of $1-1/e-\epsilon$ where $\epsilon>0$. In the following year (2010), Dolev, Elovici, and Puzis again extend the notion of group betweenness centrality to routing betweenness centrality for a traffic control network (Dolev et al., 2010).
@@
Notably, up to this point in our history, there have been no studies to focus on specific structures like cliques or walks (see Section 2), nor do they propose optimization models. One of the first OR studies in this domain is conducted and presented in Vogiatzis et al. (2015). Specifically, the study investigates the most and least important cliques in networks based on three group centrality metrics introduced in Section 2: degree, closeness, betweenness. The innovation of the work lies in the development of integer linear programming models for each variant in a compact manner. To achieve that, the authors impose an extra requirement of the size of the cliques that are being studied: this requirement has since been dropped from more recent works. Additionally, contrary to much of the more recent literature, the authors propose three variants for betweenness centrality, viewed as optimistic, pessimistic, and probabilistic. Assume that there exist more than one geodesic paths connecting two nodes, and at least one of them uses some node in the selected clique. Then, the optimistic view of betweenness assumes that this is the only shortest path used; the pessimistic view of betweenness assumes that this is not the shortest path used and another one, which is not using a node in the clique is used; finally, the probabilistic version is the traditional view of betweenness.
@@
Previous studies until 2016, primarily focus on group betweenness (with the exception of clique centrality in 2015). At this point though, the focus switches towards closeness centrality. First, Chen et al. (2016) examines group closeness centrality in order to identify a set of $k$ nodes with the highest group closeness score. The authors first propose a greedy heuristic with an approximation guarantee. To handle large-scale networks, they then develop an order-based sampling algorithm. Thereafter, we observe an increase in OR research in group centrality. We note that Bergamini et al. (2018) later improves the greedy heuristic designed by Chen et al. (2016) for the same problem through developing new techniques (bit-level parallelism, submodularity improvement). We prefer not to include this study in Figure 4. It is important to mention that group closeness is highly related to another important OR topic, that of facility location in the form of the $p$-median problem (Hakimi, 1965) as pointed out in Bergamini et al. (2018).
@@
In Veremyev et al. (2017), the researchers study group betweenness centrality and a number of variants including the length-scaled and the bounded-distance betweenness centralities. The authors examine both group of nodes and cohesive groups ($k$-clubs) and propose mixed-integer linear programming models to solve the problem. Next, Zhao et al. (Zhao et al., 2017) introduce H-group closeness centrality and investigate how to maximize this metric in large disk-resident graphs. This group centrality metric aims to identify a fixed sized node group with a distance restriction of H, which is the combination of both group degree and group closeness concepts.
@@
Still in 2017, Ohara et al. (2017) addresses a group closeness centrality concept. Very interestingly, the application studied in the manuscript ties to the evacuation performance in a given network. Specifically, the authors are interested in identifying which links to try to design (install) in order to maximize the evacuation performance. In their work, the authors show how their problem can be reduced to a node selection problem. It is the node selection phase that is performed through group centrality computation via a heuristic algorithm with the same approximation guarantee as will be shown in later studies (see, e.g., Vogiatzis and Camur (2019)). In fact, this resemblance in approximation guarantees from the greedy algorithm comes straight from submodularity and leads to an approximation guarantee of 1 - $\frac{1}{e}$.
@@
As an extension to the advancements in Vogiatzis et al. (2015) and Veremyev et al. (2017), the study conducted in Rysz et al. (2018) focuses on determining the most betweenness central clique in a connected network. The important contribution here is that, in contrast to previous works where the clique size is predefined by the authors, the cardinality restriction is lifted in Rysz et al. (2018). As is common in the literature up to now, they first develop an integer programming model. At a next step, to speed up the optimization phase, the authors then design and employ a combinatorial branch-and-bound (CBB) method. Indeed, following their work, we see a significant uptick in the number of studies utilizing CBB in group centrality literature. This fact shows that smart construction approaches can have quite the place in the group centrality literature.
@@
Interest in clique and clique relaxation problems and their intersection with centrality only continues to grow. Pastukhov et al. (2018) studies the problem of identifying a $\gamma$-quasi-clique with the highest degree. The authors first present a linear integer programming model, and then devise two exact solution approaches: a) a CBB algorithm, and b) a degree decomposition algorithm.
@@
So far, we are observing a smaller number of OR researchers and practitioners working with group centrality with a focus on optimization. This is about to change in 2019, when we witness the first surge in novel problem definitions within the OR community. Li et al. (2019) expands the concept of flow closeness centrality, proposed by Brandes and Fleischer (2005), for groups of nodes. This study carries extra significance as it includes current flow in paths other than the shortest paths in the calculation of closeness. Theoretically, the authors also derive the complexity of their problem, and propose two approximation algorithms based on the greedy algorithm. We also refer the reader to Bubboloni and Gori (2022) where the authors utilize the concepts of paths and flows to introduce new group centrality measures. However, they do not provide an optimization model or solution approaches, nor do they reference any work presented in Figure 4 that was conducted after 2006.
@@
Continuing with closeness, and similar to Ohara et al. (2017), Veremyev et al. (2019) explores a new problem in edge detection and closeness centrality. Specifically, they focus on the criticality measured by the influence of an edge’s removal on the closeness centrality of nodes connected through these edges. While Ohara et al. (2017) concentrates on new edge additions to the network, Veremyev et al. (2019) studies the implications of edge removals in relation to group closeness centrality. We note that both node and edge removal, in terms of network element criticality, are broad research topics that we refrain from delving into more in our study. Interested readers are pointed to the immense literature of network element criticality and its applications (see, e.g., Walteros and Pardalos (2012); Lalou et al. (2018)).
@@
Following up, Vogiatzis and Camur (2019) proposes a novel group centrality metric named Star Degree Centrality (SDC) where the goal is to identify the induced star with the maximum degree in a given network. The authors propose an optimization model and two approximation algorithms to demonstrate that SDC can effectively detect essential proteins in protein-protein interaction networks. The significance of this work is twofold. Firstly, it is one of the first OR studies applied in a large-scale, real-world network (motivated by protein-protein interaction networks). Their extensive computational results reveal how group centrality can play quite an important role in real-life networks, especially in contexts where nodal metrics fail. Secondly, it investigates a new cohesive group (i.e., star) following the study of cliques and their variants. Stars have a vital/critical element in their center, which makes them very unique structures in a number of applications. In a different, concurrent work, Fushimi et al. (2019) proposes connectedness centrality under stochastic edge failure, which bears similarities to group closeness centrality. The authors apply this new metric to road networks.
@@
Continuing with 2020, Rahim Taleqani et al. (2020) also focuses on maximizing closeness centrality for a connected group of nodes with a restricted diameter, also called a $k$-club (see Section 2), a similar setting observed in Veremyev et al. (2017) for betweenness. A novelty is in the application: dockless bike-sharing systems allow users to “park” their bikes anywhere, which makes the $k$-club of highest closeness particularly desirable for both users (“close” to every node in the network) and bike-sharing providers (restricted diameter to collect and find bikes). This study stands out as the first to apply closeness centrality to $k$-clubs. As a side note, there is an extensive research (see, e.g., Shahinpour and Butenko (2013)) on identifying maximum cardinality of a $k$-club (which, recall, is a clique relaxation and reduces to the clique problem for $k=1$).
We note another work titled “Group Degree Centrality and Centralization in Networks” by Krnc and Škrekovski (2020), excluded in Figure 4. The study examines set of nodes with the largest degree. However, the focus on largest-degree subgraphs without any connectivity restrictions aligns more with the maximum coverage problem (see Megiddo et al. (1983)) which is beyond our current scope. Similarly, the concept of group coverage centrality, mostly studied by researchers outside the field of OR and out of the scope the work, are not included in our review (Yoshida, 2014; Medya et al., 2018; D’Angelo et al., 2019; Komusiewicz et al., 2023), despite their importance.
@@
Within the same year, Angriman et al. (2020) proposes a new group centrality called GED-Walk centrality. This metric is motivated by Katz centrality (Katz, 1953): as such, it focuses on walks of all lengths instead of solely the shortest paths, offering an alternative to closeness and betweenness centrality measures. Following this, Akgün and Tural (2020) introduces an extension of group betweenness centrality called $k$-step centrality, where shortest path lengths are restricted by $k$. It is important to note that Zhao et al. (2017) also places length restrictions on shortest paths in their study of group closeness centrality, yet they are not cited by Akgün and Tural (2020). The methodology used by Akgün and Tural (2020) incorporates algorithms developed by Brandes (2008) and Puzis et al. (2007). In another study, Li et al. (2020) proposes a new group centrality metric named h-index group centrality, which incorporate the concept of h-index (see Hirsch (2005)) into groups.
@@
Still discussing works in 2020, we also find another group centrality study for cliques. Nasirian et al. (2020) expands the most-closeness clique problem proposed by Vogiatzis et al. (2015) into the concepts of maximum-distance-closeness centrality and total-distance closeness centrality. They reference Zhao et al. (2017) with the claim, “The maximum H-group closeness centrality problem is to detect a set of vertices of size at most H with the maximum group closeness centrality in the network”. However, we do not think that this claim is accurate since H-group closeness centrality is concerned with a combination of group degree and group closeness centrality with H being a constraint on the lengths of the shortest paths. The study conducted by Nasirian et al. (2020) shows similarity to Rysz et al. (2018), as they also remove the restriction of a fixed-sized clique, but instead focus on the most closeness-central clique, as opposed to the most betweenness-central clique.
@@
As discussed earlier, Vogiatzis et al. (2015) proposes three variants of most influential cliques with respect to degree, closeness, and betweenness. While clique betweenness is studied by both Veremyev et al. (2017) and Rysz et al. (2018), the topic of clique closeness is covered by Nasirian et al. (2020). The missing piece is completed by Zhong et al. (2021) where the authors study the most degree-central clique problem and relax the constraint on the maximum allowed size of the clique.

### v3

@@ 3 Brief history on group centrality
Figure 4: Emergence of OR scientists in group centrality metrics studies over the last two decades.
Our criteria for inclusion are as follows: the work must focus on an established or proposed centrality metric, extended and applied to a group of nodes, and solved using common OR techniques. For example, the work by Rysz et al. [83] satisfies all three criteria: the focus is on betweenness centrality, extended to groups of nodes forming cliques, and the solution technique proposed is a traditional OR technique for mathematical programs in the form of combinatorial branch-and-bound. On the other hand, the fantastic contribution by Medya et al. [61] is not included in our survey, despite its use of group betweenness as an objective function and its solution techniques being in the realm of OR. The reason is that the primary focus of the work is not on group centrality, but instead on network design. Specifically, their goal is to improve on a group centrality metric through the careful addition of edges in the network.
3.1 Foundational Studies in Group Centrality
@@
Pioneering the field are Everett and Borgatti [29] who are the first in the OR literature to discuss extending centrality and importance to groups of nodes in a network. They motivate the question for groups that are known a priori. We quote from their work: “Are the lawyers more central than the accountants in a given organisation’s social network? Is one particular ethnic minority more integrated into the community than another? To what extent are particular groups or classes (women, the elderly, African-Americans, etc.) marginalized in different networks?”. This a priori availability of group information that can help us derive descriptive information about the group itself. However, as the authors also note in that same article, we may also focus on identifying groups based on the structure of the network itself [29]. They do mention cliques as a motivation for clique centrality; see the later work by Vogiatzis et al. [102].
@@
In their original [29] and subsequent work [30], they base their proposed metrics to nodal centrality metrics. Specifically, they extend traditional metrics in the form of degree, closeness, betweenness, and flow centrality to accommodate a group setting. The authors and subsequent studies have shown how individual nodal calculations for centrality do not necessarily capture calculations that are necessary for group centrality. In the early 2000s, a fundamental study by Newman and Girvan [71] (not included in Figure 4) uses betweenness centrality to identify groups within a network. They do so by calculating edge betweenness and removing the edge of highest score until the network is disconnected. Importantly, this metric is recalculated after every removal. As this work has an arguably different focus than our review, we do not include it in our history. For completeness, though, and as readers may want to investigate the intersection between community detection (see also 3.6) and centrality, we offer the interesting comparison made by Danon et al. in [23].
@@
Continuing with these earlier years of group importance, Borgatti [13] utilizes group centrality in social network analysis to solve a key player problem. The key player problem (also known as the critical node detection problem) tasks itself with identifying a set of nodes that are “critical” for the well-being or success of a network. More recent extensions have since focused on identifying sets of network elements (nodes, edges, cliques). Borgatti [13] proposes one of the very first combinatorial optimization approaches in this domain. To our best knowledge, this is one of the earliest studies where optimization algorithms (in the form of genetic algorithms) are employed in the context of group centrality. The importance of players are measured through the impact of their absence on the graph cohesiveness. The author also discusses how degree, betweenness, and closeness metrics may fall short when applied “as is” for this problem definition of key players. Subsequently, in [77], Puzis et al. study the problem of determining groups with the largest betweenness in a large-scale network. They apply their algorithms on a network service provider infrastructure data set to optimally distribute intrusion detection devices.
@@
Another prominent work by Brandes [15] presents a wide range of variants of betweenness centrality including proximal betweenness, bounded-distance betweenness, distance-scaled betweenness, and edge betweenness. That same work also presents an algorithm to compute the group betweenness of a “given" group. It is critical to mention that the paper focuses on computing the centrality of a fixed group. This is an important step, and one that subsequent studies have used. In this work, the high complexity associated with identifying the optimal group in terms of group betweenness is also mentioned as an important factor. Thus, the study lacks algorithmic details in detecting the most betweenness group, different than Puzis et al. [77].
@@
Nonetheless, Dolev et al. [27] further this research by studying group betweenness centrality measured as the information flow in an evolving network. The authors propose a heuristic approach with an approximation guarantee that generalizes the heuristic algorithm proposed in Puzis et al. [77] for a similar problem setting. Later, Fink and Spoerhase [31] improve the analysis of Dolev et al. [27]. However, since none of these algorithms scale well in complex networks, in [59], the authors design a randomized algorithm guaranteeing an approximation of $1-1/e-\epsilon$ where $\epsilon>0$. In the following year, Dolev et al. [26] again extend the notion of group betweenness centrality to routing betweenness centrality for a traffic control network. Notably, up to this point in our history, there have been no studies to focus on specific structures like cliques or walks (see Section 2), nor do they propose optimization models.
3.2 Early Advancements and Optimization Models
@@
As the field matured, subsequent studies increasingly focused on more sophisticated group centrality metrics, incorporating combinatorial optimization techniques. One of the first OR studies in this domain is conducted and presented by Vogiatzis et al. [102]. Specifically, the study investigates the most and least important cliques in networks based on three group centrality metrics introduced in Section 2: degree, closeness, betweenness. The innovation of the work lies in the development of integer linear programming models for each variant in a compact manner. To achieve that, the authors impose an extra requirement of the size of the cliques that are being studied: this requirement has since been dropped from more recent works. Additionally, contrary to much of the more recent literature, the authors propose three variants for betweenness centrality, viewed as optimistic, pessimistic, and probabilistic. Assume that there exist more than one geodesic paths connecting two nodes, and at least one of them uses some node in the selected clique. Then, the optimistic view of betweenness assumes that this is the only shortest path used; the pessimistic view of betweenness assumes that this is not the shortest path used and another one, which is not using a node in the clique is used; finally, the probabilistic version is the traditional view of betweenness.
@@
Previous studies until 2016 primarily focus on group betweenness (with the exception of clique centrality in 2015). At this point though, the focus switches towards closeness centrality. First, Chen et al. [22] examine group closeness centrality in order to identify a set of $k$ nodes with the highest group closeness score. The authors first propose a greedy heuristic with an approximation guarantee. To handle large-scale networks, they then develop an order-based sampling algorithm. Thereafter, we observe an increase in OR research in group centrality. We note that Bergamini et al. [10] later improve the greedy heuristic designed by Chen et al. [22] for the same problem through developing new techniques (bit-level parallelism, submodularity improvement). We prefer not to include this study in Figure 4. It is important to mention that group closeness is highly related to another important OR topic, that of facility location in the form of the $p$-median problem [44] (see subsection 3.6).
@@
Veremyev et al. [97] study group betweenness centrality and a number of variants including the length-scaled and the bounded-distance betweenness centralities. The authors examine both group of nodes and cohesive groups ($k$-clubs) and propose mixed-integer linear programming models to solve the problem. Next, Zhao et al. [109] introduce H-group closeness centrality and investigate how to maximize this metric in large disk-resident graphs. This group centrality metric aims to identify a fixed sized node group with a distance restriction of H, which is the combination of both group degree and group closeness concepts.
@@
Still in 2017, Ohara et al. [73] address a group closeness centrality concept. The application studied in the manuscript ties to evacuation performance in a given network. Specifically, the authors are interested in identifying which links to try to design (install) in order to maximize the evacuation performance. In their work, the authors show how their problem can be reduced to a node selection problem. It is the node selection phase that is performed through group centrality computation via a heuristic algorithm with the same approximation guarantee as will be shown in later studies (see, e.g., [99]). In fact, this resemblance in approximation guarantees from the greedy algorithm comes straight from submodularity and leads to an approximation guarantee of 1 - $\frac{1}{e}$.
3.3 Emergence of Optimization Algorithms
@@
The evolution of group centrality has been marked by significant advancements in optimization models and algorithmic strategies. As an extension to the advancements in Vogiatzis et al. [102] and Veremyev et al. [97], the study conducted by Rysz et al. [83] focuses on determining the most betweenness central clique in a connected network. The important contribution here is that, in contrast to previous works where the clique size is predefined by the authors, the cardinality restriction is lifted. As is common in the literature up to now, they first develop an integer programming model. At a next step, to speed up the optimization phase, the authors then design and employ a combinatorial branch-and-bound (CBB) method. Indeed, following their work, we see a significant uptick in the number of studies utilizing CBB in group centrality literature. This fact shows that smart construction approaches can have quite the place in the group centrality literature.
@@
Interest in clique and clique relaxation problems and their intersection with centrality only continues to grow. Pastukhov et al. [76] study the problem of identifying a $\gamma$-quasi-clique with the highest degree. The authors first present a linear integer programming model, and then devise two exact solution approaches: a) a CBB algorithm, and b) a degree decomposition algorithm.
@@
So far, we are observing a smaller number of OR researchers working with group centrality with a focus on optimization. This is about to change in 2019, when we witness the first surge in novel problem definitions within the OR community. Li et al. [56] expand the concept of flow closeness centrality, proposed by Brandes and Fleischer [16], for groups of nodes. This study carries extra significance as it includes current flow in paths other than the shortest paths in the calculation of closeness. Theoretically, the authors also derive the complexity of their problem, and propose two approximation algorithms based on the greedy algorithm. We also refer the reader to Bubboloni and Gori [18] where the authors utilize the concepts of paths and flows to introduce new group centrality measures. However, they do not provide an optimization model or solution approaches, nor do they reference any work presented in Figure 4 that was conducted after 2006.
@@
Continuing with closeness, and similar to Ohara et al. [73], Veremyev et al. [98] explore a new problem in edge detection and closeness centrality. Specifically, they focus on the criticality measured by the influence of an edge’s removal on the closeness centrality of nodes connected through these edges. While Ohara et al. [73] concentrates on new edge additions to the network, Veremyev et al. [98] study the implications of edge removals in relation to group closeness centrality. We note that both node and edge removal, in terms of network element criticality, are broad research topics that we refrain from delving into more in our study. Interested readers are pointed to the immense literature of network element criticality and its applications (see, e.g., [104, 55]). We mention these more concisely in subsection 3.6.
3.4 Trends and Novel Group Centrality Metrics
@@
Emerging trends in group centrality research have introduced novel concepts and metrics, such as probabilistic models and new structural definitions, reflecting the continuous innovation within the field. Following up, Vogiatzis and Camur [99] propose a novel group centrality metric named Star Degree Centrality (SDC) where the goal is to identify the induced star with the maximum degree in a given network. The authors propose an optimization model and two approximation algorithms to demonstrate that SDC can effectively detect essential proteins in protein-protein interaction networks. The significance of this work is twofold. Firstly, it is one of the first OR studies applied in a large-scale, real-world network (motivated by protein-protein interaction networks). Their extensive computational results reveal how group centrality can play quite an important role in real-life networks, especially in contexts where nodal metrics fail. Secondly, it investigates a new cohesive group (i.e., star) following the study of cliques and their variants. Stars have a vital/critical element in their center, which makes them very unique structures in a number of applications. In a different, concurrent work, Fushimi et al. [40] propose connectedness centrality under stochastic edge failure, which bears similarities to group closeness centrality. The authors apply this new metric to road networks.
@@
Continuing with 2020, Taleqani et al. [78] focus on maximizing closeness centrality for a connected group of nodes with a restricted diameter, also called a $k$-club (see Section 2), a similar setting observed in Veremyev et al. [97] for betweenness. A novelty is in the application: dockless bike-sharing systems allow users to “park” their bikes anywhere, which makes the $k$-club of highest closeness particularly desirable for both users (“close” to every node in the network) and bike-sharing providers (restricted diameter to collect and find bikes). This study stands out as the first to apply closeness centrality to $k$-clubs. As a side note, there is an extensive research (see, e.g., [87]) on identifying maximum cardinality of a $k$-club (which, recall, is a clique relaxation and reduces to the clique problem for $k=1$).
@@
Within the same year, Angriman et al. [4] propose a new group centrality called GED-Walk centrality. This metric is motivated by Katz centrality [50]: as such, it focuses on walks of all lengths instead of solely the shortest paths, offering an alternative to closeness and betweenness centrality measures. Following this, Akgün and Tural [2] introduce an extension of group betweenness centrality called $k$-step centrality, where shortest path lengths are restricted by $k$. It is important to note that Zhao et al. [109] also place length restrictions on shortest paths in their study of group closeness centrality, yet they are not cited by Akgün and Tural [2]. The methodology used by Akgün and Tural [2] incorporate algorithms developed by Brandes [15] and Puzis et al. [77]. In another study, Li et al.  [57] propose a new group centrality metric named h-index group centrality, which incorporate the concept of h-index (see [47]) into groups.
@@
Still discussing works in 2020, we also find another group centrality study for cliques. Nasirian et al. [68] expand the most-closeness clique problem proposed by Vogiatzis et al. [102] into the concepts of maximum-distance-closeness centrality and total-distance closeness centrality. They reference Zhao et al. [109] with the claim, “The maximum H-group closeness centrality problem is to detect a set of vertices of size at most H with the maximum group closeness centrality in the network". However, we do not think that this claim is accurate since H-group closeness centrality is concerned with a combination of group degree and group closeness centrality with H being a constraint on the lengths of the shortest paths. The study conducted by Nasirian et al. [68] shows similarity to Rysz et al. [83], as they also remove the restriction of a fixed-sized clique, but instead focus on the most closeness-central clique, as opposed to the most betweenness-central clique.
@@
As discussed earlier, Vogiatzis et al. [102] propose three variants of most influential cliques with respect to degree, closeness, and betweenness. While clique betweenness is studied by both Veremyev et al. [97] and Rysz et al. [83], the topic of clique closeness is covered by Nasirian et al. [68]. The missing piece is completed by Zhong et al. [110] where the authors study the most degree-central clique problem and relax the constraint on the maximum allowed size of the clique.
3.5 Recent Developments in Group Centrality